OpenStudy (anonymous):

Can someone please help? I need to know the difference between recursive and explicit equations

OpenStudy (anonymous):

The sum of the first "n" numbers (= 1+2+ 3 + ...+ n) Explicit means that you can get the answer directly from the formula; Example: S(n) = (1/2)(n)(n+1) Recursive means that you can get the next number from the current one Example: S(n+1) = S(n) + (n+1)

OpenStudy (anonymous):

Thankyou so much . I understand this better now